OpenStudy (anonymous):

During World War I, the Germans had a gun called Big Bertha that was used to shell Paris. The shell had an initial speed of 0.765 km/s at an initial inclination of 36.2◦ to the horizontal. The acceleration of gravity is 9.8 m/s^2. How far away did the shell hit? Answer in units of km

OpenStudy (agent0smith):

Are you familiar with the range equation? \[\huge R = \frac{ v_o ^2 \sin (2 \theta) }{ g }\] if so you can use it here. If not you'll have to start by breaking the velocity up into horiz and vertical components.

OpenStudy (anonymous):

I'm not familiar with the range equation either...

OpenStudy (agent0smith):

Start by finding the components of v then  use sine and cosine to find Vx and Vy.

OpenStudy (anonymous):

I got 569.22 but it's not right.

OpenStudy (anonymous):

I mean I got 569.22 for the final answer...

OpenStudy (agent0smith):

Well idk what the final answer is, but i can work it out with the range equation... but for now just work out Vx and Vy

OpenStudy (agent0smith):

I put V in m/s just to be safe, we can get km/s back later

OpenStudy (anonymous):

So Vy is 451.81 and Vx is 617.32?

OpenStudy (agent0smith):

Very good :D Now, find the time taken, using Vy. When the object hits the ground, Vy will be equal but in the opposite direction... ie -451.81 So use the equation \[\Large v_f = v_i + a t\] where vi is 451.81, vf is -451.81, and a=-9.8m/s^2 (negative since it's in the same direction as the downward v)... find t.

OpenStudy (agent0smith):

Once you have t, just use distance = Vx*time btw it looks like your answer was only off by a factor of 10... 569.22 when it should be more like 56.92 km.

OpenStudy (anonymous):

Oh yeah... it is 56.92! I can't believe I messed up at the easiest part of the problem haha. ^-^ thanks.

OpenStudy (agent0smith):

haha i figured i'd check and thought you'd just forgotten to divide by g somewhere! :)

OpenStudy (anonymous):

Well I got a 98% on my homework so I'm pretty happy. (:

OpenStudy (agent0smith):

aw yay! :D